the length of a rectangle is 8 meters less than 5 times its width.if the perimeter of the rectangle is at most 104 meters, find the greatest possible length of the rectangle
Answer by checkley71(8403)

L=5W-8
PERIMETER=2L+2W
104=2(5W-8)+2W
104=10W-16+2W
12W=104=16
W=120/12
W=10 ANSWER FOR THE WIDTH.
L=5*10-8
L=42 ANSWER.
PROOF:
104=2*42+2*10
104=84+20
104=104